### 03 Mar 2026 v1

# arXiv:2603.03152v1[econ.GN] 03 Mar 2026

Abstract
Using transaction-level matched trades from Polymarket’s 2024 U.S.presidential-election contracts, we study how prediction markets process major political shocks. We focus on three events with precise timestamps: the first Biden-Trump debate, the Trump assassination attempt, and Biden’s drop out. We document large bursts of activity on both extensive and intensive margins, concentrated among high-intensity incumbents, and show that pre-event net exposure predicts abnormal post-event trading and position flips. To link order flow to prices, we estimate a Kyle-style price-impact measure and a Glosten–Harris decomposition that separates permanent from transitory order-flow effects, complemented by variance-ratio dynamics and a bounded two-sidedness index. Across shocks, price discovery differs sharply: the debate exhibits stronger transitory pressure and partial reversal, the assassination attempt features a more permanent repricing, and the drop out episode combines heavy trading with muted net price changes and high two-sidedness, consistent with disagreement under Knightian uncertainty.
Keywords: prediction markets; price discovery; political shocks; market liquidity; trader heterogeneity; Knightian uncertainty; Kyle’s $\lambda$.
JEL Codes: G14; D82; D84; G41.
1 Introduction
Prediction markets are often interpreted as real-time probabilities of future political outcomes, so whether they incorporate major news quickly and efficiently matters for researchers, forecasters, and policymakers. Yet even when news is public and highly salient, transaction prices can move sharply and then drift or reverse, raising a basic question: when a political shock hits, do prices primarily reflect belief updating, or do they partly reflect transitory price pressure generated by order flow in a thin market?
In this study, we examine Polymarket’s 2024 U.S. presidential-election contracts using transaction-level matched trades recorded on-chain. Our empirical setting centers on three political shocks with precise timestamps: the first Biden-Trump debate (2024-06-28, 1:00 UTC), the assassination attempt on Donald Trump (2024-07-13, 22:11 UTC), and President Biden’s drop out from the race (2024-07-21, 17:46 UTC). These events span anticipated versus unprecedented shocks and allow us to measure both trading and price adjustment in 5-minute event time.
We first document how traders respond to these shocks. Trading rises sharply on both margins: new address entry spikes after the assassination attempt and the drop out announcement, and incumbent traders’ activity increases markedly after each event. The marginal response is highly concentrated among high-intensity traders, and pre-event net exposure to the “Trump wins” state predicts both abnormal post-event trading and subsequent position flips, consistent with active risk management rather than passive rebalancing.
We then ask how this event-time trading translates into prices. Table 1 illustrates the heterogeneity in adjustment paths for the Trump YES token: the debate produces a sharp move followed by partial reversal, the assassination attempt yields a large and persistent repricing, and the drop out episode shows heavy trading with a modest net price change but a sizeable interim drawdown. These patterns motivate our mechanism analysis of price discovery under limited liquidity and disagreement.
Our analysis proceeds in two steps. We first document the event-time trading response and identify which trader types supply the marginal flow. We then study event-time price discovery using three complementary lenses: (i) reduced-form price impact and a permanent–transitory decomposition of order-flow effects, (ii) dynamic diagnostics based on variance ratios that distinguish drift from reversal, and (iii) a bounded two-sidedness measure that captures how balanced buy and sell pressure is when the market processes ambiguous news. This structure lets us separate short-lived price pressure from more durable belief updates and interpret why some shocks generate intense trading with muted net price changes.
We contribute to the prediction-market and microstructure literatures by providing high-frequency evidence on trading and participation around major political shocks in a large, decentralized market; by linking order flow to prices through a Kyle-style price-impact measure and a GH decomposition tailored to binary contracts; and by diagnosing post-event price dynamics and disagreement using variance ratios and a bounded two-sidedness index that connects naturally to Knightian uncertainty.
The remainder of the paper proceeds as follows. Section 2 describes Polymarket’s institutional details, the data construction, and our measurement choices. Section 3 and Section 4 study event-time trading responses and their concentration across traders. Section 5subsection 5.3 analyze event-time price adjustment and mechanisms. subsection 5.4 discusses interpretation and open questions, and Section 6 concludes.
1.1 Related Literature
Prediction markets trade state-contingent claims whose prices can be interpreted as aggregate beliefs about future events. A large literature studies their forecasting performance and discusses design and regulatory issues (Wolfers and Zitzewitz, 2004; Arrow et al., 2008; Snowberg et al., 2013). Political prediction markets, one category of prediction markets, have received growing scholarly and public attention. Rhode and Strumpf (2004) analyze large, well-organized U.S. presidential betting markets operating between 1868 and 1940 and show that the odds forecasted elections remarkably well and were fairly efficient despite active attempts to manipulate them. Work on the Iowa Electronic Markets shows that election prediction markets are at least as accurate as, and often more accurate than, opinion polls, especially at longer horizons (Berg et al., 2008a, b). A related concern is susceptibility to manipulation in thin markets: Rhode and Strumpf (2007) study speculative attacks in historical Wall Street markets, the Iowa Electronic Markets, and TradeSports, and find price impacts that are quickly undone with little evidence of systematic manipulation beyond short horizons. We contribute to this literature by using a large, high-frequency dataset from Polymarket to characterize how trading activity and liquidity respond to salient political shocks (Baker et al., 2016; Rothschild and Sethi, 2016).
Our analysis also relates to the market-microstructure literature on price impact and the role of order flow in price discovery. Canonical models such as Kyle (1985) and Glosten and Milgrom (1985) imply that the mapping from signed order flow to prices summarizes how information is incorporated into transaction prices. Empirical microstructure work measures how trades move prices and develops econometric tools to decompose price changes into informed and uninformed components (Hasbrouck, 1991, 1995; Easley et al., 1997). We adapt standard microstructure measurement—event-time activity analysis and a Kyle-style price-impact proxy—to a binary prediction-market setting.
Finally, the paper speaks to evidence on heterogeneity across traders and the role of experience and attention. Prior work using brokerage and fintech data documents that trading behavior varies sharply across individuals and over time (Barber and Odean, 2000; Seru et al., 2010; Barber et al., 2022). In the prediction-market context, we focus on a simple and economically meaningful dimension of heterogeneity: whether traders were already active prior to major news and how intensively they traded before the shock.
2 Institutional Background and Data
2.1 Institutional background: Polymarket market design
Polymarket operates as a decentralized prediction market on the Polygon blockchain, using the Gnosis Conditional Tokens Framework. Trades settle on-chain via smart contracts while order matching occurs through a centralized limit order book. All transactions use USDC stablecoin. The Presidential Election Winner 2024 market lists 17 candidates as binary outcome tokens; we focus on Trump, Biden, and Harris, which account for 95% of trading volume.
The Gnosis framework supports three transaction types: simple trades (secondary market exchange), minting (creating YES/NO pairs by locking \$1 USDC collateral), and burning (redeeming paired tokens). Minting and burning are primarily arbitrage and liquidity-provision activities rather than directional bets, a distinction critical for interpreting volume data.
2.2 Data and construction
2.2.1 Data collection
Each matched trade on Polymarket generates an immutable on-chain record that includes traded value in USDC, traded token quantity, timestamp, and the addresses of both counterparties. These records are emitted as logs by the exchange smart contracts, providing a complete transaction-level history suitable for studying price discovery and liquidity in a high-frequency setting. Our empirical application focuses on the Presidential Election Winner 2024 market, which is implemented as a multi-outcome market, and the corresponding logs are emitted by the smart contract NegRisk CTFExchange. We collect the relevant on-chain logs directly from the Polygon blockchain via the Polygon RPC endpoint.
We restrict attention to January 5, 2024 (market launch) through November 6, 2024 (after Fox News first projected the election winner). The Presidential Election Winner 2024 market lists 17 candidates, however, here we focus on the three candidates that account for the majority of trading activity in our sample: Donald Trump, Joe Biden, and Kamala Harris. Over this period, we collect 3,654,974 matched trades involving 232,469 unique addresses for these three candidates.
2.2.2 Filtering systematic and advanced operator addresses
A small number of addresses appear to act not merely based on their own belief about the final outcome, but also based on platform mechanics. We identify such addresses using the following heuristic criteria:
  1. 1.
    Known Polymarket addresses. If an address is known to be controlled by Polymarket, we label it as such.
  2. 2.
    Position conversion activity. Addresses that conduct Position Conversion via the smart contract NegRisk Adapter are flagged as advanced operators, as this activity is commonly associated with arbitrage or inventory management by advanced users.
  3. 3.
    Negative token holdings. Using our reconstructed token holding dataset from exchange-trade logs, we flag addresses whose token holdings have ever turned negative for any candidate, which typically indicates off-exchange inventory changes or conversion activity not captured in the trade logs emitted by NegRisk CTFExchange.
Negative token holdings can arise for two reasons. First, position conversions executed via the NegRisk Adapter do not appear in the NegRisk CTFExchange trade logs used to reconstruct our token holdings dataset. Hence, the subsequent sales of converted tokens would appear as sales of inventory “never purchased” in our dataset. Second, addresses may transfer tokens across wallets (e.g., from a cold wallet to a hot wallet) that occur outside the exchange and being unobservable in our trading dataset. Then, subsequent exchange trades can again create negative holdings as if the address never owned the tokens. In our data, only seven addresses satisfy the latter criterion without also exhibiting conversion activity. Out of an abundance of caution, we exclude all of them from our analysis.
Finally, we remove clusters of addresses that transfer tokens among themselves outside the exchange. These addresses are highly likely controlled by the same underlying entity for operational or security reasons. Applying these filters yields 2,176 advanced operator addresses and 355 addresses that are likely controlled by the same entities.
Our transaction data provide executed trades but not unexecuted order book depth, motivating transaction-based liquidity measures rather than quoted spreads. We infer trade direction using standard tick-rule methods. These constraints are common in blockchain-based market analysis and do not preclude testing theoretical predictions about adjustment dynamics.
2.3 Event design and measurement map
As mentioned in Section 1, we focus on three political shocks with well-defined timestamps (Biden-Trump Debate, Trump assassination attempt, and Biden’s drop out). Throughout the paper we aggregate trades into 5-minute bins (ceiling). Each examined event has also been assigned to a 5-minute bin for consistency in event-time indexing.
Our transaction data include executed trades but not full order-book depth, motivating us to adopt transaction-based liquidity and price-impact measures rather than quoted spreads. When trade direction is needed, we infer buyer-initiated versus seller-initiated trades using a tick-rule classifier (Hasbrouck, 1991).
Different sections use different windows tailored to the object of interest: the trading-related plots and trader heterogeneity regressions use a 30-minute pre/post window with trader characteristics measured in the month before the event. We choose this time window to isolate the effect of the event from the following up events. In later price-discovery diagnostics, we use a common 2-hour pre and 4-hour post window to fully illustrate the price discovery pattern around each event. Table 2 summarizes the main variables used in this paper and their measure frequency.
3 Event-Driven Trading Activity
This section documents how Polymarket’s presidential-election markets react in event time to the three political shocks in Section 1. We summarize market-level adjustment around these events and then focus on 5-minute event time to characterize (i) new address entry (the extensive margin) and (ii) abnormal trading by incumbent traders (the intensive margin).
3.1 Overall market activity
Figure 1 displays the evolution of prices and transaction volume in the prediction markets for Trump, Biden, and Harris over the period spanning these three events.
Over this period, the prices of the YES and NO tokens move gradually but exhibit several sharp revaluations that coincide with major events. The price of the Biden YES token begins to drop sharply after the first presidential debate. Around the same time, the price of the Harris YES token begins to move away from its near-zero level, roughly 23 days before President Biden officially withdrew from the race and endorsed Vice President Harris as the Democratic nominee. The Harris YES token price then rises rapidly to around 0.4 after 2024-07-21, when President Biden officially endorsed her. The joint movement in (Trump YES, Harris YES) (and, equivalently, (Trump NO, Harris NO)) is consistent with a reallocation of market-implied probability weight toward President Trump and Vice President Harris as the two main candidates for the 2024 presidential election.
Based on the transaction-volume data, the Trump market dominates over this period. Transaction volume in the Trump market is much higher than in the other two markets before 2024-07-21, and transaction volume in the Harris market begins to catch up after the announcement of President Biden’s drop out.
3.2 Extensive margin: new traders’ entry
In Figure 2 and Figure 3, we examine how traders react to these three events at a 5-minute frequency. Figure 2 plots the number of new traders entering the presidential election markets in each 5-minute interval around the three major political events. There are substantial spikes in new entry around the Trump assassination attempt and Biden’s drop out, indicating that these shocks attract large numbers of new participants into the presidential markets. For the first presidential debate, the number of new traders is relatively low, which may reflect two factors: (1) at the time of the debate, Polymarket was still in an earlier stage of adoption; and (2) the debate, while salient, was less sudden and dramatic than the other two events, which are rare in the context of modern U.S. presidential elections.
3.2.1 Intensive margin: incumbent traders’ trading activity
For incumbent traders, following Brown and Warner (1985) and MacKinlay (1997), we implement a standard event-study design at a 5-minute resolution. For each event $e$, we re-index trading into event-time bins around $t=0$ and measure trader-level outcomes including trading volume and trading frequency. Meanwhile, because a single limit order can be split into multiple transactions and potentially inflate measured trading frequency, we also employ trading intensity, which is defined as the number of 5-minute intervals in which a trader is active, as a robustness check. We construct abnormal trading by comparing event-window outcomes to a trader-specific pre-event baseline computed in an event-specific estimation window (the three hours before each event), using an inverse-hyperbolic-sine transformation to accommodate zeros while preserving a log-like scale. We then aggregate abnormal outcomes across traders within each event-time bin and report cumulative abnormal activity with 95% confidence intervals. Appendix A provides full definitions and implementation details.
[…]
These results show that shocks generate large bursts of trading on both the extensive and intensive margins. We next ask who supplies this marginal event-time order flow, because composition and concentration affect how order flow maps into prices.
4 Heterogeneity and Concentration of Trading Responses
As shown in Section 3, trading activity rises sharply around the three political shocks we study, reflecting both newcomer entry and increased activity among traders already present on the platform. This section asks which traders drive the response. We focus on incumbent traders and relate abnormal post-event trading in a 30-minute window to (i) historical trading intensity (volume and frequency) and (ii) pre-event net exposure to the “Trump wins” state. Additional characteristics, including single-market participation and simple contrarian/momentum indicators, are included as secondary controls.
4.1 Trader heterogeneity and position management
Consistent with the above setting, we still divide time into 5-minute intervals and focus on the event window $(t-30\text{ minutes},\,t+30\text{ minutes})$ [^1] We choose a 30-minute window to reduce contamination from closely timed follow-up events. For example: (1) the assassination attempt on President Trump occurred at 10:11 PM UTC, and Elon Musk publicly endorsed Trump at 10:45 PM UTC; (2) President Biden dropped out of the race at 5:46 PM UTC, and he announced his support for Vice President Harris at 6:13 PM UTC.. Meanwhile, we use trading data from $(t-1\text{ month},\,t-30\text{ minutes})$ as an estimation window to construct trader characteristics. These characteristics fall into four dimensions: (1) trading volume and trading frequency, (2) exposure to net Trump WIN tokens, (3) single-market participation, and (4) contrarian versus momentum behavior. We classify a trader as contrarian/momentum in a given token market if, over the sample period, the trader’s signed trading value is negatively/positively correlated with the lagged 60-minute rolling price trend.
Because traders can hold both YES and NO tokens for each candidate, we construct each trader’s net Trump WIN token position by combining positions across the three main candidates (Trump, Biden, and Harris):
$\text{Net Trump Win}=$ $\text{Trump YES}+\text{Biden NO}+\text{Harris NO}$
$-\text{Trump NO}-\text{Biden YES}-\text{Harris YES}$
We then divide traders into two groups based on their net Trump WIN position: (1) positive net Trump WIN position and (2) negative net Trump WIN position [^2] Fewer than 50 traders have a zero net Trump WIN position in the estimation window. Because this is negligible relative to the total number of traders, we group these traders with the positive-net-position group.. These constructed trader characteristics and corresponding summary statistics are summarized in Table 3 and Table 4.
We next estimate regressions to examine how trader characteristics relate to trading activity around the three events. First, we pool the three events and estimate a panel regression with trader and time fixed effects:
$y_{it}=\alpha+\bm{\beta_{1}}\bm{X_{it}}\times\text{Post Event}_{t}+\epsilon_{it}$ (1)
where $y_{it}$ is trader $i$’s abnormal trading behavior at time $t$, $\text{Post Event}_{t}$ is an indicator equal to 1 after the event time, and $\bm{X_{it}}$ is the vector of trader characteristics defined in Table 3. We report standard errors clustered at the trader level to account for within-trader dependence over time and across events.
Table 5 and Table 6 shows high-volume and high-frequency traders remain disproportionately active post-event, suggesting trading concentration. More revealing: traders with negative Trump WIN exposure (who profit if Trump loses) become significantly more active after all three events, consistent with active position management. This exposure effect is economically meaningful: traders positioned against Trump increase abnormal volume by 0.034 log-units (column 4), roughly matching the baseline post-event surge.
One concern with trading frequency is that a single order may be split into multiple trades, potentially inflating measured frequency. To address this concern, we also use trading intensity as an alternative dependent variable. The trading intensity is measured by how many unique 5-minute intervals the trader is active in. Table C.1 reports similar patterns.
We also estimate pooled OLS regressions of the form:
$y_{it}=\alpha+\beta_{1}\text{Post Event}_{t}+\bm{\beta_{2}}\bm{X_{it}}+\bm{\beta_{3}}\bm{X_{it}}\times\text{Post Event}_{t}+\epsilon_{it}$ (2)
where $y_{it}$ is trader $i$’s abnormal trading behavior at time $t$, $\text{Post Event}_{t}$ is an indicator equal to 1 after the event time, and $\bm{X_{it}}$ is the vector of trader characteristics defined in Table 3.
Table C.2 and Table C.3 report pooled OLS results for trading volume and trading frequency. The results are consistent with the panel regressions. In particular, traders with negative net Trump WIN exposure are not always more active on average, but they become significantly more active than traders with positive exposure in the post-event window.
To reduce repetition in the main text, we report event-specific panel regressions in Appendix D. The event-by-event estimates show that the sign of exposure-related activity can differ across shocks. For the debate (when the Trump WIN price moves up) and Biden’s drop out (when it moves down then have a quick rebound), traders with negative net Trump WIN exposure exhibit larger abnormal activity in the immediate post-event 30 minutes window. For the assassination attempt, the exposure pattern reverses over the first 30 minutes. These patterns are consistent with short-run position adjustment around shocks. We also estimate event-specific pooled OLS regressions and obtain qualitatively similar conclusions (see Appendix D).
4.2 Position flips and short-run position adjustment
We next examine whether elevated post-event trading coincides with changes in net exposure by studying position flips. We construct a binary indicator for whether a trader flips their net Trump WIN position after the event (i.e., moving from a negative position to zero/positive, or from a nonnegative position to negative). We estimate the following model:
$PF_{it}=\alpha+\bm{\beta}\bm{X_{it}}+\epsilon_{it}$ (3)
where $PF_{it}$ is the indicator for a post-event position flip and $\bm{X_{it}}$ is the vector of trader characteristics defined in Table 3. Because the dependent variable is binary, we estimate a generalized linear model (GLM) with a binomial family and logit link, which directly models $P\left(Y_{i,e,t}=1\mid X_{i,e,t}\right)$ and ensures fitted probabilities lie in $[0,1]$. We report standard errors clustered at the trader level to account for within-trader dependence over time and across events.
Table 7 reports results for position flips. In the pooled specification (column (1)), traders with negative net Trump WIN exposure are more likely to flip their position after the event. When we estimate the model separately by event, the pattern differs across events. For the debate and Biden dropping out, negative-exposure traders are more likely to flip (columns (2) and (4)). For the assassination-attempt event, positive-exposure traders are more likely to flip (column (3)). Overall, these results are consistent with elevated post-event trading being accompanied by active position adjustment around shocks.
5 Event-Time Price Discovery and Liquidity
Section 3Section 4 show that political shocks generate large bursts of trading and that the marginal order flow is concentrated in specific trader groups. We now ask how this event-time trading translates into prices. Because the on-chain record reveals executed trades but not order-book depth, we focus on transaction prices and volumes aggregated to 5-minute event time in a common [-2h, +4h] window around each shock.
5.1 Stylized facts: price and volume around shocks
Figure 4 plots transaction prices and trading volume for Trump YES/NO tokens around the three events. The shapes differ across events: the debate features a sharp move followed by partial reversal; the assassination attempt features a jump and a more persistent re-leveling; and the Biden drop out event features comparatively muted price changes with a large volume response. Appendix E reports analogous plots for Biden and Harris tokens.
We interpret these adjustment patterns using three complementary lenses: (i) reduced-form price impact and a permanent–transitory decomposition of order-flow effects, (ii) dynamic diagnostics that assess whether post-event evolution reflects drift or reversal, and (iii) a bounded two-sidedness measure that captures how balanced buy and sell pressure is when the market processes ambiguous news.
5.2 Order-flow price impact: Kyle’s $\lambda$ and the GH decomposition
In a thin market, transaction prices can move even in the absence of new information because trades consume limited liquidity. A Kyle-style regression provides a summary measure of how sensitive log-odds prices are to net signed order flow; we interpret a larger $\lambda$ as lower effective depth. However, $\lambda$ alone does not distinguish information-driven belief updating from temporary price pressure. We therefore complement it with the Glosten–Harris (GH) decomposition, which separates a more permanent component linked to the level of signed flow from a more transitory component linked to changes in signed flow.
Among the outcome tokens we examine, the Trump YES token attracts the largest trading volume and provides the richest transaction-level record in both coverage and time span. We therefore estimate Kyle’s $\lambda$ and the Glosten–Harris decomposition for the Trump YES market using detailed trade data. Our approach follows standard high-frequency microstructure practice, adapted to the setting of a binary prediction market, and the full procedure is detailed in Appendix B.
Figure 5 summarizes the event-time estimates. Panel (a) reports $\lambda$, the slope of log-odds changes on net signed order flow in 5-minute bins. Importantly, because $\lambda$ is estimated on net flow, it can appear elevated when trading is highly two-sided and net flow varies little relative to gross volume. As we show in subsection 5.4, this is especially relevant for the Biden drop out episode, which features intense but largely offsetting buy and sell pressure and therefore a modest net price change despite large volume. A matched weekday-time placebo test of the same statistic yields p=0.086 (debate), p=0.002 (assassination), and p=0.022 (drop out), supporting that the event-time change in price impact is more than a generic same-time pattern (see Figure C.2).
Panel (b) presents the key decomposition result. The coefficient on $Q_{t}$ is interpreted as the more permanent (information-related) component, while the coefficient on $\Delta Q_{t}$ reflects a more transitory component tied to temporary price pressure (inventory/spread effects). In the debate, the transitory component is relatively large compared to the permanent component, consistent with a sharp move followed by partial reversal. In the assassination episode, the permanent component dominates and the post-event price level shift largely persists. In the drop out episode, both components are sizable but the net price change is small, consistent with balanced two-sided trading and small realized net order flow. Overall, the decomposition indicates that event-time price jumps combine belief updating with liquidity-driven pressure, and the persistence of those jumps depends on the relative importance of the permanent versus transitory component.
5.3 Post-event dynamics: reversal vs. drift
If the decomposition indicates a large transitory component, we should observe short-horizon reversals as liquidity pressure unwinds. This section assesses whether post-event price evolution reflects efficient price discovery or short-run drift/reversal using rolling variance ratios computed from 5-minute log-odds returns with a 6-bin horizon (30 minutes) (Lo and MacKinlay, 1988). The variance ratio is defined as $VR(6)=\frac{Var(r_{t}^{(30\text{ min})})}{6Var(r_{t}^{(5\text{ min})})}$ and $r_{t}$ represents the log-odds change, so that under a random-walk benchmark for efficient price discovery, $VR(6)\approx 1$. Values $VR(6)>1$ indicate short-horizon drift (positive serial correlation), while $VR(6)<1$ indicates short-horizon mean reversion (negative serial correlation).
Figure 6 shows that event-time dynamics differ across shocks. The debate exhibits pronounced mean reversion after the initial move, consistent with the relatively large transitory component in the GH decomposition. The assassination episode displays a short-lived drift phase ($VR(6)>1$) during the immediate repricing and then reverts toward values closer to one, consistent with a rapid transition toward a new level. The Biden drop out episode remains well below one, consistent with persistent short-horizon reversals and the two-sided, liquidity-driven trading pressure highlighted below. Because we use transaction prices rather than midquotes, $VR(6)<1$ can also mechanically reflect microstructure frictions (bid–ask bounce, discreteness, and transient order-flow pressure). We therefore interpret the variance ratio as a diagnostic of drift versus reversal rather than a literal test of market efficiency. Using a matched weekday-time placebo test based on the post-event maximum of $VR(6)$ in the first post-event hour, the assassination episode shows an unusually large drift spike (max $VR(6)\approx 1.84$, p=0.002), the debate is marginal (max $\approx 1.28$, p=0.078), and the drop out episode shows no such spike (max $\approx 0.49$, p=0.675), as expected given its below-one VR pattern (see Figure C.3).
5.4 Knightian uncertainty: two-sided trading pressure
A final ingredient is ambiguity (Knightian uncertainty): the same headline can map into election probabilities in multiple plausible ways, so traders may disagree about the correct post-event price even when the information set is public (Knight, 1921; Epstein and Wang, 1994; Illeditsch, 2011; Pástor and Veronesi, 2012, 2013). In a prediction market, such disagreement can generate heavy trading with limited net order flow and therefore muted net price changes.
To capture this “two-sided” trading in a way that is well behaved even when price changes are small, we compute a bounded order-imbalance index using buy and sell volume within each 5-minute bin, $\text{TwoSided}_{t}\equiv 1-\frac{|B_{t}-S_{t}|}{B_{t}+S_{t}},$ where $B_{t}$ and $S_{t}$ are buy- and sell-initiated volume (USDC). The index lies in $[0,1]$: values near one indicate balanced buy/sell pressure (high two-sidedness), while values near zero indicate one-sided flow.
Figure 7 shows that two-sidedness differs sharply across shocks. The Biden drop out episode features the highest two-sidedness around the event, consistent with a large volume response but muted net price changes. In contrast, the debate is initially one-sided and becomes progressively more balanced in the following hours, consistent with the partial reversal pattern. The assassination attempt starts from a one-sided repricing phase and then transitions toward more balanced trading as the market settles. A matched weekday-time placebo test shows the two-sidedness measure is insignificant for the debate (p=0.523), consistent with the absence of a discrete immediate jump, but significant for the assassination attempt (p=0.014) and the drop out episode (p=0.002), consistent with a shift toward more balanced trading pressure (see Figure C.4).
Taken together, the bounded two-sidedness index helps reconcile why we sometimes observe intense trading without a commensurate price move and why Kyle’s $\lambda$ (estimated on net flow) can increase even when gross volume is high. Combined with the GH decomposition and the variance-ratio diagnostics, it suggests that price discovery under political shocks reflects an interaction between liquidity, heterogeneity in beliefs, and the persistence of order flow.
6 Conclusion
This paper uses transaction-level matched trades from Polymarket’s 2024 U.S. presidential-election contracts to study how prediction markets process major political shocks. Using event time around the first Biden-Trump debate, the Trump assassination attempt, and President Biden’s drop out, we document sharp increases in participation and trading volume. Entry spikes are especially large after the two unprecedented shocks, and the marginal response is concentrated among high-intensity incumbent traders. Pre-event net exposure to the “Trump wins” state predicts abnormal post-event trading and position flips, consistent with active position management around shocks.
We then connect order flow to prices. A Kyle-style price-impact measure changes around shocks, and the Glosten–Harris decomposition shows that the composition of price impact differs across events: the debate is associated with relatively stronger transitory pressure and partial reversal, the assassination attempt with a more permanent repricing, and the drop out episode with intense but largely two-sided trading and muted net price changes. Variance-ratio dynamics provide complementary evidence on drift versus reversal, and the bounded two-sidedness index summarizes the degree of balanced trading pressure consistent with disagreement under Knightian uncertainty.
Our analysis highlights both the promise and the limits of prediction markets as real-time probabilities during high-salience events: prices can revalue rapidly, but transaction-level microstructure and heterogeneous beliefs shape the path and persistence of adjustment. A key limitation of on-chain data is that we observe executed trades but not unexecuted order-book depth, so our liquidity measures are necessarily transaction-based. Future work combining on-chain transactions with order-book snapshots and cross-market linkages could further separate belief updating from liquidity provision and risk-transfer motives.
Appendix Online Appendix
This appendix contains five items. Appendix A details the event-study construction used in the main text. Appendix B details the construction of Kyle’s $\lambda$ and the Glosten–Harris specification in the binary prediction-market setting. Appendix C reports robustness checks and supplementary specifications, including the matched weekday-time placebo tests. Appendix D reports event-specific panel regressions that are omitted from the main text for brevity. Appendix E reports additional event-time price adjustment plots referenced in the main text.
Appendix A Event-Study Method
This method follows Brown and Warner (1985) and MacKinlay (1997). We conduct an event-study analysis of trading responses around these three events. For each event $e$ occurring at time $t=0$, we reorganize the data into 5-minute event time, $k\in\mathbb{Z}$, where $k$ indexes 5-minute bins relative to the event timestamp.
For each trader $i$, event $e$, and event-time bin $k$, we measure trading outcomes on the intensive margin: trading volume $V_{i,e,k}$, trading frequency $F_{i,e,k}$, and an indicator for participation $D_{i,e,k}=\mathbbm{1}[V_{i,e,k}>0]$. To accommodate zeros while retaining an approximately log-like scale for large values, we apply the inverse hyperbolic sine transformation $g(x)=\operatorname{asinh}(x)$ to volume and frequency.
(1) Trader-specific baseline (estimation window)
To construct abnormal trading, we estimate each trader’s normal trading intensity using an event-specific estimation window prior to the event. Let $\mathcal{K}^{\text{est}}_{e}$ denote the set of 5-minute bins in the pre-event estimation window. For each trader $i$, we compute baseline trading intensity as the within-trader mean during $\mathcal{K}^{\text{est}}_{e}$:
$\bar{V}_{i,e}$ $=\frac{1}{|\mathcal{K}^{\text{est}}_{e}|}\sum_{k\in\mathcal{K}^{\text{est}}_{e}}V_{i,e,k},$ $\bar{F}_{i,e}$ $=\frac{1}{|\mathcal{K}^{\text{est}}_{e}|}\sum_{k\in\mathcal{K}^{\text{est}}_{e}}F_{i,e,k},$ (4)
$\bar{D}_{i,e}$ $=\frac{1}{|\mathcal{K}^{\text{est}}_{e}|}\sum_{k\in\mathcal{K}^{\text{est}}_{e}}\mathbbm{1}[V_{i,e,k}>0].$ (5)
We then form transformed baselines $g(\bar{V}_{i,e})$ and $g(\bar{F}_{i,e})$.
(2) Abnormal trading in the event window
Let $\mathcal{K}^{\text{evt}}_{e}$ denote the event window around $t=0$. We define abnormal trading for each trader and bin as deviations from the trader-specific baseline:
$AV_{i,e,k}$ $=g(V_{i,e,k})-g(\bar{V}_{i,e}),$ (6)
$AF_{i,e,k}$ $=g(F_{i,e,k})-g(\bar{F}_{i,e}),$ (7)
$AP_{i,e,k}$ $=D_{i,e,k}-\bar{D}_{i,e}.$ (8)
For volume, we additionally set $AV_{i,e,k}=0$ when $V_{i,e,k}=0$, so that bins with no trading are treated as having no abnormal activity on the transformed scale.
(3) Aggregation and cumulative abnormal activity
We aggregate abnormal outcomes across traders within each event-time bin. For outcome $Y\in\{AV,AF,AP\}$, the mean abnormal outcome at event time $k$ is:
$\widehat{\mu}^{Y}_{e,k}=\frac{1}{N_{e,k}}\sum_{i}Y_{i,e,k},$ (9)
where $N_{e,k}$ is the number of trader-bin observations available at $(e,k)$. To summarize total post-event responses, we compute cumulative abnormal activity (CAA) as the cumulative sum of bin-level mean abnormal outcomes:
$CAA^{Y}_{e,\tau}=\sum_{k\leq\tau}\widehat{\mu}^{Y}_{e,k}.$ (10)
We report $95\%$ confidence intervals for $CAA^{Y}_{e,\tau}$ using standard errors computed from the cross-sectional dispersion of $Y_{i,e,k}$ within each bin and propagating uncertainty through the cumulative sum under the simplifying assumption of independence across bins:
$\widehat{se}\!\left(CAA^{Y}_{e,\tau}\right)$ $=\sqrt{\sum_{k\leq\tau}\left(\widehat{\sigma}^{Y}_{e,k}/\sqrt{N_{e,k}}\right)^{2}},$ (11)
$CAA^{Y}_{e\tau}$ $\pm 1.96\cdot\widehat{se}\!\left(CAA^{Y}_{e,\tau}\right),$ (12)
where $\widehat{\sigma}^{Y}_{e,k}$ is the cross-sectional standard deviation of $Y_{i,e,k}$ at event time $k$.
Appendix B Kyle’s $\lambda$
This section details how Kyle’s $\lambda$ is adopted to the setting of a binary prediction market. We estimate Kyle’s $\lambda$ for the Trump YES market using detailed trade data.
(1) Infer the direction of transactions
Because we do not observe the full limit order book and only observe matched trades recorded on the Polygon blockchain, we infer trade direction (buyer- versus seller-initiated). We apply a tick-rule classifier: we assign direction $d_{n}=\operatorname{sign}(p_{n}-p_{n-1})\in\{+1,-1\}$, and when the price does not change we carry forward the most recent non-zero direction (Hasbrouck, 1991). We then define signed order flow as:
$q_{n}=d_{n}\cdot tx_{n},$ (13)
where $tx_{n}$ is the transaction size in terms of million USDs. This signed volume $q_{n}$ captures the net demand pressure each transaction places on the market.
(2) Time aggregation
Prediction-market trading is irregular, so to map order flow to price impact consistently we aggregate trades into fixed 5-minute bins in event time. For each bin $t$, we compute the volume-weighted average price (VWAP) $p_{t}^{\text{VWAP}}$ and net order flow $Q_{t}$ as:
$p_{t}^{\text{VWAP}}=\frac{\sum_{n\in t}w_{n}p_{n}}{\sum_{n\in t}w_{n}},\qquad Q_{t}=\sum_{n\in t}q_{n},$ (14)
If no trades occurred in a given bin, we carry forward the last price and set $Q_{t}=0$ to maintain a continuous time series.
(3) Log-Odds Price Transformation
Because a prediction market price is essentially a probability bounded between 0 and 1, price changes are naturally heteroskedastic (e.g. a 5-cent move is more significant when the price is \$0.1 vs. when it’s \$0.5). To account for this, we convert prices to log-odds to obtain an unbounded metric that is more comparable across different price levels and mitigate nonlinearity and heteroskedasticity near 0 and 1 (Wolfers and Zitzewitz, 2006):
$\theta_{t}=\ln\left(\frac{p_{t}}{1-p_{t}}\right).$ (15)
We then take first differences $\Delta\theta_{t}=\theta_{t}-\theta_{t-1}$ as our measure of price movement in each 5-minute bin, analogous to a return.
(4) Event-time estimation of price impact
We estimate Kyle’s $\lambda$ on 5-minute bins within each event window by regressing log-odds changes on signed order flow:
$\Delta\theta_{t}=\lambda\,Q_{t}+\varepsilon_{t}$ (16)
To separate permanent from transitory effects, we estimate the Glosten–Harris specification:
$\Delta\theta_{t}=\alpha+\lambda_{\text{perm}}Q_{t}+\lambda_{\text{trans}}\Delta Q_{t}+\varepsilon_{t}$ (17)
with HAC standard errors. As an estimate of price impact, $\lambda$ is a reduced-form liquidity proxy: larger values indicate lower effective depth, but it does not distinguish informed trading from noise trading. We therefore interpret $\lambda$ alongside the permanent/transitory decomposition in event time.
Appendix C Robustness Checks
C.1 Supplementary regression results
Table C.1: Panel Regression Results (5-min interval), Trading Intensity as the Y
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0039***
Post Event $\times$ Trad Vol High 0.0110*** 0.0061***
(17.199) (12.150)
Post Event $\times$ Trad Int Multi 0.0145*** 0.0103***
(15.511) (11.256)
Post Event $\times$ Neg Trump Win 0.0066*** 0.0042***
(11.418) (8.2673)
Post Event $\times$ Single Mkt -0.0021** 0.0007
(-2.0800) (0.6706)
Post Event $\times$ Contrarian 0.0049** -0.0046*
(2.0365) (-1.8407)
Post Event $\times$ Momentum 0.0138*** 0.0038
(3.1258) (0.8513)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0017 0.0046 0.0053 0.0024 -0.0020 0.0006 0.0063
R-Squared (Between) 0.0207 0.0495 0.0575 0.0240 -0.0189 0.0059 0.0710
R-Squared (Overall) 0.0040 0.0100 0.0116 0.0050 -0.0041 0.0012 0.0141
Observations 284420 284420 284420 284420 284420 284420 284420
  • Notes: This table repeats the stacked panel design in Table 5 and Table 6, using abnormal trading intensity (participation) as the dependent variable.t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table C.2: OLS Regression Results (5-min interval), Trading Volume as the Y
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0231*** 0.0019*** 0.0063*** 0.0126*** 0.0324*** 0.0215*** -0.0141**
(0.0013) (0.0004) (0.0008) (0.0012) (0.0059) (0.0013) (0.0061)
Trad Vol High 0.0037*** 0.0029***
(0.0011) (0.0007)
Post Event $\times$ Trad Vol High 0.0685*** 0.0453***
(0.0041) (0.0036)
Trad Freq High 0.0039*** 0.0008
(0.0014) (0.0013)
Post Event $\times$ Trad Freq High 0.0703*** 0.0412***
(0.0048) (0.0047)
Neg Trump Win 0.0014 0.0006
(0.0009) (0.0009)
Post Event $\times$ Neg Trump Win 0.0339*** 0.0218***
(0.0034) (0.0032)
Single Mkt -0.0028** -0.0025
(0.0013) (0.0016)
Post Event $\times$ Single Mkt -0.0103* 0.0069
(0.0061) (0.0064)
Contrarian 0.0150 0.0130
(0.0092) (0.0092)
Post Event $\times$ Contrarian 0.0324** -0.0113
(0.0137) (0.0143)
Momentum 0.0064** 0.0041
(0.0031) (0.0033)
Post Event $\times$ Momentum 0.0762** 0.0308
(0.0303) (0.0302)
R-squared 0.0018 0.0083 0.0077 0.0034 0.0019 0.0026 0.0107
R-squared Adj. 0.0018 0.0083 0.0076 0.0034 0.0019 0.0026 0.0106
Observations 293,624 293,624 293,624 293,624 293,624 293,624 293,624
  • Notes: Unit of observation is a trader-by-5-minute bin in the event window, pooled across the three events. The dependent variable is trader-level abnormal trading volume in event time. Post Event equals 1 for bins at and after the event time. Each column reports an OLS specification with event fixed effects. Regressors include Post Event, the indicated pre-event trader characteristic (Table 3), and its interaction with Post Event. Standard errors are clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table C.3: OLS Regression Results (5-min interval), Trading Frequency as the Y
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0042*** 0.0006*** 0.0011*** 0.0020*** 0.0070*** 0.0039*** -0.0017
(0.0003) (0.0001) (0.0001) (0.0002) (0.0014) (0.0002) (0.0013)
Trad Vol High 0.0007*** 0.0006***
(0.0002) (0.0001)
Post Event $\times$ Trad Vol High 0.0116*** 0.0070***
(0.0008) (0.0007)
Trad Freq High 0.0007** 0.0001
(0.0003) (0.0003)
Post Event $\times$ Trad Freq High 0.0129*** 0.0080***
(0.0009) (0.0009)
Neg Trump Win 0.0002 0.0001
(0.0002) (0.0002)
Post Event $\times$ Neg Trump Win 0.0070*** 0.0048***
(0.0007) (0.0006)
Single Mkt -0.0008*** -0.0008**
(0.0003) (0.0004)
Post Event $\times$ Single Mkt -0.0032** 0.0003
(0.0014) (0.0014)
Contrarian 0.0025 0.0021
(0.0016) (0.0016)
Post Event $\times$ Contrarian 0.0050** -0.0029
(0.0024) (0.0025)
Momentum 0.0010** 0.0006
(0.0005) (0.0005)
Post Event $\times$ Momentum 0.0149** 0.0067
(0.0068) (0.0068)
R-squared 0.0018 0.0075 0.0077 0.0038 0.0021 0.0026 0.0105
R-squared Adj. 0.0018 0.0075 0.0077 0.0038 0.0021 0.0026 0.0104
Observations 293,624 293,624 293,624 293,624 293,624 293,624 293,624
  • Notes: This table is a repetition of Table C.2, using abnormal trading frequency as the dependent variable. Standard errors are clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table C.4: OLS Regression Results (5-min interval), Trading Intensity as the Y
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0040*** 0.0006*** 0.0011*** 0.0019*** 0.0061*** 0.0037*** -0.0016
(0.0002) (0.0001) (0.0001) (0.0002) (0.0011) (0.0002) (0.0010)
Trad Vol High 0.0007*** 0.0004***
(0.0002) (0.0001)
Post Event $\times$ Trad Vol High 0.0109*** 0.0061***
(0.0007) (0.0005)
Neg Trump Win 0.0003 0.0000
(0.0002) (0.0002)
Post Event $\times$ Neg Trump Win 0.0065*** 0.0042***
(0.0006) (0.0005)
Single Mkt -0.0009*** -0.0008**
(0.0003) (0.0003)
Post Event $\times$ Single Mkt -0.0024** 0.0004
(0.0011) (0.0011)
Contrarian 0.0024 0.0018
(0.0015) (0.0016)
Post Event $\times$ Contrarian 0.0049** -0.0046**
(0.0021) (0.0022)
Momentum 0.0010* 0.0003
(0.0005) (0.0006)
Post Event $\times$ Momentum 0.0137** 0.0038
(0.0059) (0.0059)
Post Event $\times$ Trad Int Multi 0.0144*** 0.0103***
(0.0010) (0.0009)
Trad Int Multi 0.0010*** 0.0005
(0.0003) (0.0003)
R-squared 0.0020 0.0082 0.0101 0.0041 0.0023 0.0029 0.0126
R-squared Adj. 0.0020 0.0082 0.0101 0.0041 0.0022 0.0029 0.0126
Observations 293,624 293,624 293,624 293,624 293,624 293,624 293,624
  • Notes: This table is a repetition of Table C.2, using abnormal trading intensity (participation) as the dependent variable. Standard errors are clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
C.2 Matched weekday-time placebo tests
[…]
(a) Abnormal incumbent trading volume
(b) Abnormal incumbent trading frequency
(c) Abnormal incumbent trading intensity
Figure C.1: Placebo distributions for intensive-margin abnormal activity
Notes: Each histogram shows the placebo distribution of the event-time jump in abnormal activity, computed at 500 pseudo-event timestamps matched on weekday and clock time; the dashed vertical line marks the real event value. Panels are arranged as a 3x3 grid with columns for debate/assassination/drop out and rows for volume/frequency/intensity. Randomization p-values for these events are: volume 0.012/0.032/0.002; frequency 0.030/0.018/0.002; intensity 0.030/0.002/0.002.
(a) Biden-Trump debate
(b) Trump assassination
(c) Biden drop out
Figure C.2: Placebo distributions for Kyle’s $\lambda$
Notes: Each histogram shows the placebo distribution of the event-time change in Kyle’s $\lambda$, computed at 500 pseudo-event timestamps matched on weekday and clock time; the dashed vertical line marks the real event value. Randomization p-values are: debate 0.086, assassination 0.002, drop out 0.022.
(a) Biden-Trump debate
(b) Trump assassination
(c) Biden drop out
Figure C.3: Placebo distributions for variance-ratio spikes (VR(6))
Notes: Each histogram shows the placebo distribution of the post-event maximum of $VR(6)$ in the first post-event hour, computed at 500 pseudo-event timestamps matched on weekday and clock time; the dashed vertical line marks the real event value. Real-event maxima and p-values are: debate max $\approx 1.28$ (p=0.078), assassination max $\approx 1.84$ (p=0.002), drop out max $\approx 0.49$ (p=0.675).
(a) Biden-Trump debate
(b) Trump assassination
(c) Biden drop out
Figure C.4: Placebo distributions for two-sidedness
Notes: Each histogram shows the placebo distribution of the event-time increase in the two-sidedness index, computed at 500 pseudo-event timestamps matched on weekday and clock time; the dashed vertical line marks the real event value. Randomization p-values are: debate 0.523, assassination 0.014, drop out 0.002.
Appendix D Event-Specific Panel Regressions
Table D.1: Panel Regression Results (5-min interval): Biden-Trump Debate (Trading Volume)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0049***
Post Event $\times$ Trad Vol High 0.0109*** 0.0077**
(3.0511) (2.3758)
Post Event $\times$ Trad Freq High 0.0100** -0.0043
(2.2981) (-1.1288)
Post Event $\times$ Neg Trump Win 0.0189*** 0.0182***
(5.3808) (5.3966)
Post Event $\times$ Single Mkt -0.0052 -0.0039
(-0.7408) (-0.5118)
Post Event $\times$ Contrarian 0.0088 0.0092
(0.5270) (0.5424)
Post Event $\times$ Momentum 0.0421** 0.0433**
(2.4691) (2.4410)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0003 0.0005 0.0004 0.0011 -0.0007 0.0008 0.0019
R-Squared (Between) 0.0059 0.0090 0.0068 0.0114 -0.0086 0.0091 0.0204
R-Squared (Overall) 0.0009 0.0015 0.0011 0.0023 -0.0017 0.0018 0.0041
Observations 71581 71581 71581 71581 71581 71581 71581
  • Notes: This table reports event-specific panel regressions following the same setting in Table 5. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.2: Panel Regression Results (5-min interval): Biden-Trump Debate (Trading Frequency)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0011***
Post Event $\times$ Trad Vol High 0.0026*** 0.0016**
(3.1609) (2.1682)
Post Event $\times$ Trad Freq High 0.0030*** 7.124e-05
(2.8300) (0.0779)
Post Event $\times$ Neg Trump Win 0.0041*** 0.0039***
(5.1558) (5.6354)
Post Event $\times$ Single Mkt -0.0015 -0.0010
(-0.7071) (-0.4333)
Post Event $\times$ Contrarian 0.0029 0.0022
(0.7970) (0.5869)
Post Event $\times$ Momentum 0.0089** 0.0083**
(2.3971) (2.1123)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0003 0.0006 0.0006 0.0012 -0.0011 0.0009 0.0020
R-Squared (Between) 0.0058 0.0086 0.0087 0.0106 -0.0114 0.0092 0.0201
R-Squared (Overall) 0.0009 0.0016 0.0016 0.0024 -0.0024 0.0019 0.0043
Observations 71581 71581 71581 71581 71581 71581 71581
  • Notes: This table reports event-specific panel regressions following the same setting in Table 6. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.3: Panel Regression Results (5-min interval): Biden-Trump Debate (Trading Intensity)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0009***
Post Event $\times$ Trad Vol High 0.0022*** 0.0012**
(3.2910) (2.1141)
Post Event $\times$ Trad Int Multi 0.0030*** 0.0001
(2.9441) (0.1327)
Post Event $\times$ Neg Trump Win 0.0038*** 0.0037***
(5.7039) (5.7823)
Post Event $\times$ Single Mkt -0.0003 0.0003
(-0.1962) (0.1686)
Post Event $\times$ Contrarian 0.0030 0.0023
(0.9283) (0.7045)
Post Event $\times$ Momentum 0.0080*** 0.0075**
(2.5934) (2.3204)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0003 0.0006 0.0006 0.0012 -0.0002 0.0009 0.0017
R-Squared (Between) 0.0065 0.0089 0.0100 0.0118 -0.0021 0.0104 0.0203
R-Squared (Overall) 0.0010 0.0016 0.0018 0.0025 -0.0004 0.0020 0.0039
Observations 71581 71581 71581 71581 71581 71581 71581
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.1. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.4: Panel Regression Results (5-min interval): Trump Assassination (Trading Volume)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0253***
Post Event $\times$ Trad Vol High 0.0814*** 0.0697***
(10.412) (8.5666)
Post Event $\times$ Trad Freq High 0.0682*** 0.0238**
(7.4440) (2.3953)
Post Event $\times$ Neg Trump Win -0.0101** -0.0231***
(-2.0884) (-4.5700)
Post Event $\times$ Single Mkt -0.0069 -0.0033
(-0.7464) (-0.3278)
Post Event $\times$ Contrarian 0.1477*** 0.1002**
(3.5134) (2.4352)
Post Event $\times$ Momentum 0.1083** 0.0645
(2.0360) (1.2238)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0017 0.0055 0.0035 -0.0004 -0.0009 0.0014 0.0063
R-Squared (Between) 0.0192 0.0572 0.0370 -0.0043 -0.0094 0.0152 0.0657
R-Squared (Overall) 0.0036 0.0111 0.0072 -0.0008 -0.0019 0.0029 0.0128
Observations 93200 93200 93200 93200 93200 93200 93200
  • Notes: This table reports event-specific panel regressions following the same setting in Table 5. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.5: Panel Regression Results (5-min interval): Trump Assassination (Trading Frequency)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0040***
Post Event $\times$ Trad Vol High 0.0124*** 0.0098***
(9.6797) (7.8691)
Post Event $\times$ Trad Freq High 0.0118*** 0.0056***
(7.3552) (3.2898)
Post Event $\times$ Neg Trump Win -0.0012 -0.0034***
(-1.4461) (-3.8601)
Post Event $\times$ Single Mkt -0.0016 -0.0004
(-0.9156) (-0.2374)
Post Event $\times$ Contrarian 0.0204*** 0.0122**
(3.3625) (2.0416)
Post Event $\times$ Momentum 0.0195* 0.0119
(1.8103) (1.1103)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0016 0.0048 0.0037 -0.0003 -0.0013 0.0012 0.0057
R-Squared (Between) 0.0183 0.0498 0.0385 -0.0031 -0.0126 0.0126 0.0588
R-Squared (Overall) 0.0035 0.0098 0.0076 -0.0006 -0.0026 0.0025 0.0117
Observations 93200 93200 93200 93200 93200 93200 93200
  • Notes: This table reports event-specific panel regressions following the same setting in Table 6. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.6: Panel Regression Results (5-min interval): Trump Assassination (Trading Intensity)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0038***
Post Event $\times$ Trad Vol High 0.0116*** 0.0089***
(10.381) (8.2492)
Post Event $\times$ Trad Int Multi 0.0121*** 0.0061***
(7.6464) (3.7683)
Post Event $\times$ Neg Trump Win -0.0013* -0.0034***
(-1.8291) (-4.6206)
Post Event $\times$ Single Mkt -0.0011 -0.0005
(-0.7311) (-0.3654)
Post Event $\times$ Contrarian 0.0207*** 0.0124**
(3.5074) (2.1234)
Post Event $\times$ Momentum 0.0195* 0.0117
(1.8847) (1.1365)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0018 0.0051 0.0040 -0.0004 -0.0010 0.0015 0.0062
R-Squared (Between) 0.0215 0.0574 0.0463 -0.0041 -0.0105 0.0164 0.0694
R-Squared (Overall) 0.0039 0.0108 0.0086 -0.0008 -0.0020 0.0031 0.0130
Observations 93200 93200 93200 93200 93200 93200 93200
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.1. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.7: Panel Regression Results (5-min interval): Biden Drop Out (Trading Volume)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0321***
Post Event $\times$ Trad Vol High 0.0923*** 0.0479***
(13.240) (8.9342)
Post Event $\times$ Trad Freq High 0.1015*** 0.0625***
(12.808) (8.9922)
Post Event $\times$ Neg Trump Win 0.0787*** 0.0612***
(11.369) (9.8864)
Post Event $\times$ Single Mkt -0.0129 0.0251**
(-1.2900) (2.3995)
Post Event $\times$ Contrarian -0.0167 -0.0499*
(-0.6462) (-1.7529)
Post Event $\times$ Momentum 0.1711** 0.1250
(1.9649) (1.4623)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0023 0.0065 0.0069 0.0051 -0.0019 0.0007 0.0059
R-Squared (Between) 0.0252 0.0670 0.0700 0.0506 -0.0176 0.0052 0.0743
R-Squared (Overall) 0.0051 0.0141 0.0148 0.0108 -0.0039 0.0012 0.0144
Observations 119639 119639 119639 119639 119639 119639 119639
  • Notes: This table reports event-specific panel regressions following the same setting in Table 5. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.8: Panel Regression Results (5-min interval): Biden Drop Out (Trading Frequency)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0061***
Post Event $\times$ Trad Vol High 0.0162*** 0.0080***
(11.674) (6.7893)
Post Event $\times$ Trad Freq High 0.0185*** 0.0113***
(11.537) (7.6092)
Post Event $\times$ Neg Trump Win 0.0154*** 0.0121***
(10.742) (9.9064)
Post Event $\times$ Single Mkt -0.0048** 0.0024
(-1.9867) (1.0113)
Post Event $\times$ Contrarian -0.0025 -0.0084*
(-0.6706) (-1.9452)
Post Event $\times$ Momentum 0.0330* 0.0245
(1.7682) (1.3378)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0024 0.0061 0.0067 0.0055 -0.0044 0.0007 0.0078
R-Squared (Between) 0.0222 0.0537 0.0585 0.0464 -0.0337 0.0048 0.0748
R-Squared (Overall) 0.0052 0.0130 0.0142 0.0114 -0.0086 0.0013 0.0175
Observations 119639 119639 119639 119639 119639 119639 119639
  • Notes: This table reports event-specific panel regressions following the same setting in Table 6. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.9: Panel Regression Results (5-min interval): Biden Drop Out (Trading Intensity)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0058***
Post Event $\times$ Trad Vol High 0.0155*** 0.0067***
(13.504) (8.2522)
Post Event $\times$ Trad Int Multi 0.0213*** 0.0149***
(13.366) (10.667)
Post Event $\times$ Neg Trump Win 0.0146*** 0.0109***
(12.415) (10.806)
Post Event $\times$ Single Mkt -0.0038** 0.0026
(-2.0449) (1.4303)
Post Event $\times$ Contrarian -0.0040* -0.0117***
(-1.8641) (-4.1205)
Post Event $\times$ Momentum 0.0296* 0.0189
(1.9257) (1.2528)
Entity Effects YES YES YES YES YES YES YES
Time Effects NO YES YES YES YES YES YES
R-Squared (Within) 0.0026 0.0066 0.0086 0.0060 -0.0039 0.0007 0.0092
R-Squared (Between) 0.0284 0.0681 0.0879 0.0584 -0.0346 0.0053 0.1037
R-Squared (Overall) 0.0059 0.0145 0.0188 0.0127 -0.0078 0.0013 0.0213
Observations 119639 119639 119639 119639 119639 119639 119639
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.1. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.10: OLS Regression Results (5-min interval): Biden-Trump Debate (Trading Volume)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0050*** 0.0017* 0.0029*** -0.0009 0.0106* 0.0033*** 0.0008
(0.0012) (0.0009) (0.0010) (0.0009) (0.0064) (0.0010) (0.0065)
Trad Vol High 0.0024 0.0026*
(0.0020) (0.0015)
Post Event $\times$ Trad Vol High 0.0107*** 0.0077**
(0.0034) (0.0031)
Trad Freq High 0.0030 -0.0007
(0.0027) (0.0017)
Post Event $\times$ Trad Freq High 0.0099** -0.0047
(0.0042) (0.0036)
Neg Trump Win -0.0016 -0.0026*
(0.0011) (0.0014)
Post Event $\times$ Neg Trump Win 0.0188*** 0.0180***
(0.0034) (0.0032)
Single Mkt -0.0057* -0.0070**
(0.0032) (0.0036)
Post Event $\times$ Single Mkt -0.0062 -0.0051
(0.0065) (0.0070)
Contrarian 0.0149 0.0150
(0.0161) (0.0161)
Post Event $\times$ Contrarian 0.0087 0.0095
(0.0161) (0.0163)
Momentum -0.0013*** -0.0013
(0.0005) (0.0017)
Post Event $\times$ Momentum 0.0420** 0.0435**
(0.0164) (0.0170)
R-squared 0.0003 0.0011 0.0009 0.0016 0.0006 0.0019 0.0039
R-squared Adj. 0.0002 0.0010 0.0009 0.0016 0.0005 0.0018 0.0037
Observations 71,581 71,581 71,581 71,581 71,581 71,581 71,581
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.2. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.11: OLS Regression Results (5-min interval): Biden-Trump Debate (Trading Frequency)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0011*** 0.0003* 0.0005** -0.0002 0.0029 0.0007*** 0.0004
(0.0003) (0.0002) (0.0002) (0.0002) (0.0019) (0.0002) (0.0018)
Trad Vol High 0.0003 0.0005*
(0.0004) (0.0003)
Post Event $\times$ Trad Vol High 0.0025*** 0.0016**
(0.0008) (0.0007)
Trad Freq High 0.0004 -0.0004
(0.0005) (0.0003)
Post Event $\times$ Trad Freq High 0.0029*** -0.0001
(0.0010) (0.0009)
Neg Trump Win -0.0004* -0.0006**
(0.0002) (0.0003)
Post Event $\times$ Neg Trump Win 0.0041*** 0.0038***
(0.0008) (0.0006)
Single Mkt -0.0016** -0.0020**
(0.0008) (0.0009)
Post Event $\times$ Single Mkt -0.0019 -0.0015
(0.0019) (0.0020)
Contrarian 0.0027 0.0030
(0.0030) (0.0030)
Post Event $\times$ Contrarian 0.0028 0.0023
(0.0035) (0.0035)
Momentum -0.0003*** -0.0000
(0.0001) (0.0003)
Post Event $\times$ Momentum 0.0088** 0.0084**
(0.0036) (0.0038)
R-squared 0.0003 0.0011 0.0012 0.0017 0.0009 0.0019 0.0042
R-squared Adj. 0.0003 0.0010 0.0012 0.0016 0.0009 0.0018 0.0040
Observations 71,581 71,581 71,581 71,581 71,581 71,581 71,581
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.3. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.12: OLS Regression Results (5-min interval): Biden-Trump Debate (Trading Intensity)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0010*** 0.0003* 0.0005** -0.0002 0.0016 0.0006*** -0.0007
(0.0002) (0.0002) (0.0002) (0.0002) (0.0013) (0.0002) (0.0013)
Trad Vol High 0.0003 0.0003
(0.0004) (0.0002)
Post Event $\times$ Trad Vol High 0.0021*** 0.0012**
(0.0006) (0.0005)
Trad Int Multi 0.0006 0.0001
(0.0006) (0.0004)
Post Event $\times$ Trad Int Multi 0.0029*** 0.0001
(0.0010) (0.0008)
Neg Trump Win -0.0004* -0.0006**
(0.0002) (0.0003)
Post Event $\times$ Neg Trump Win 0.0037*** 0.0036***
(0.0006) (0.0006)
Single Mkt -0.0017** -0.0019**
(0.0008) (0.0008)
Post Event $\times$ Single Mkt -0.0007 -0.0001
(0.0013) (0.0014)
Contrarian 0.0025 0.0024
(0.0028) (0.0028)
Post Event $\times$ Contrarian 0.0030 0.0024
(0.0031) (0.0032)
Momentum -0.0003*** -0.0004
(0.0001) (0.0004)
Post Event $\times$ Momentum 0.0079*** 0.0075**
(0.0029) (0.0031)
R-squared 0.0003 0.0010 0.0014 0.0017 0.0007 0.0020 0.0041
R-squared Adj. 0.0003 0.0010 0.0013 0.0017 0.0006 0.0019 0.0039
Observations 71,581 71,581 71,581 71,581 71,581 71,581 71,581
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.4. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.13: OLS Regression Results (5-min interval): Trump Assassination (Trading Volume)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0254*** 0.0017*** 0.0110*** 0.0286*** 0.0323*** 0.0224*** 0.0094
(0.0023) (0.0006) (0.0016) (0.0028) (0.0084) (0.0022) (0.0090)
Trad Vol High 0.0017 0.0014*
(0.0012) (0.0008)
Post Event $\times$ Trad Vol High 0.0813*** 0.0697***
(0.0075) (0.0078)
Trad Freq High 0.0018 0.0003
(0.0015) (0.0015)
Post Event $\times$ Trad Freq High 0.0681*** 0.0236**
(0.0088) (0.0096)
Neg Trump Win 0.0003 -0.0002
(0.0008) (0.0009)
Post Event $\times$ Neg Trump Win -0.0102** -0.0233***
(0.0046) (0.0048)
Single Mkt -0.0023 -0.0021
(0.0018) (0.0021)
Post Event $\times$ Single Mkt -0.0077 -0.0042
(0.0087) (0.0093)
Contrarian -0.0007** -0.0014
(0.0004) (0.0014)
Post Event $\times$ Contrarian 0.1476*** 0.1004**
(0.0404) (0.0396)
Momentum 0.0128 0.0119
(0.0134) (0.0134)
Post Event $\times$ Momentum 0.1082** 0.0645
(0.0511) (0.0506)
R-squared 0.0016 0.0091 0.0059 0.0017 0.0016 0.0037 0.0112
R-squared Adj. 0.0016 0.0091 0.0058 0.0017 0.0016 0.0037 0.0111
Observations 93,200 93,200 93,200 93,200 93,200 93,200 93,200
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.2. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.14: OLS Regression Results (5-min interval): Trump Assassination (Trading Frequency)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0040*** 0.0004*** 0.0016*** 0.0045*** 0.0056*** 0.0036*** 0.0015
(0.0004) (0.0001) (0.0002) (0.0004) (0.0016) (0.0004) (0.0016)
Trad Vol High 0.0004 0.0003*
(0.0003) (0.0002)
Post Event $\times$ Trad Vol High 0.0124*** 0.0097***
(0.0012) (0.0012)
Trad Freq High 0.0005 0.0001
(0.0003) (0.0004)
Post Event $\times$ Trad Freq High 0.0117*** 0.0055***
(0.0015) (0.0016)
Neg Trump Win 0.0001 -0.0001
(0.0002) (0.0002)
Post Event $\times$ Neg Trump Win -0.0013 -0.0034***
(0.0008) (0.0008)
Single Mkt -0.0007 -0.0007
(0.0005) (0.0005)
Post Event $\times$ Single Mkt -0.0018 -0.0007
(0.0016) (0.0017)
Contrarian -0.0002** -0.0004
(0.0001) (0.0003)
Post Event $\times$ Contrarian 0.0203*** 0.0122**
(0.0058) (0.0058)
Momentum 0.0018 0.0015
(0.0020) (0.0020)
Post Event $\times$ Momentum 0.0194* 0.0119
(0.0103) (0.0103)
R-squared 0.0014 0.0079 0.0062 0.0015 0.0016 0.0033 0.0101
R-squared Adj. 0.0014 0.0079 0.0062 0.0015 0.0015 0.0032 0.0099
Observations 93,200 93,200 93,200 93,200 93,200 93,200 93,200
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.3. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.15: OLS Regression Results (5-min interval): Trump Assassination (Trading Intensity)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0038*** 0.0005*** 0.0017*** 0.0043*** 0.0050*** 0.0034*** 0.0017
(0.0003) (0.0001) (0.0002) (0.0004) (0.0014) (0.0003) (0.0013)
Trad Vol High 0.0005 0.0003
(0.0003) (0.0002)
Post Event $\times$ Trad Vol High 0.0116*** 0.0089***
(0.0011) (0.0010)
Trad Int Multi 0.0007 0.0004
(0.0005) (0.0005)
Post Event $\times$ Trad Int Multi 0.0121*** 0.0061***
(0.0015) (0.0016)
Neg Trump Win 0.0001 -0.0001
(0.0002) (0.0002)
Post Event $\times$ Neg Trump Win -0.0014* -0.0035***
(0.0007) (0.0007)
Single Mkt -0.0007 -0.0006
(0.0004) (0.0005)
Post Event $\times$ Single Mkt -0.0013 -0.0008
(0.0014) (0.0014)
Contrarian -0.0002** -0.0006
(0.0001) (0.0005)
Post Event $\times$ Contrarian 0.0207*** 0.0124**
(0.0057) (0.0056)
Momentum 0.0020 0.0016
(0.0022) (0.0023)
Post Event $\times$ Momentum 0.0195* 0.0117
(0.0099) (0.0099)
R-squared 0.0016 0.0086 0.0072 0.0017 0.0017 0.0039 0.0115
R-squared Adj. 0.0016 0.0086 0.0072 0.0016 0.0017 0.0038 0.0114
Observations 93,200 93,200 93,200 93,200 93,200 93,200 93,200
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.4. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.16: OLS Regression Results (5-min interval): Biden Drop Out (Trading Volume)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0322*** 0.0022*** 0.0045*** 0.0083*** 0.0447*** 0.0311*** -0.0409***
(0.0023) (0.0007) (0.0011) (0.0015) (0.0092) (0.0022) (0.0098)
Trad Vol High 0.0062*** 0.0038***
(0.0021) (0.0014)
Post Event $\times$ Trad Vol High 0.0921*** 0.0479***
(0.0067) (0.0052)
Trad Freq High 0.0066*** 0.0037*
(0.0024) (0.0022)
Post Event $\times$ Trad Freq High 0.1013*** 0.0621***
(0.0076) (0.0067)
Neg Trump Win 0.0040** 0.0027
(0.0020) (0.0019)
Post Event $\times$ Neg Trump Win 0.0785*** 0.0609***
(0.0066) (0.0059)
Single Mkt -0.0019 -0.0000
(0.0020) (0.0027)
Post Event $\times$ Single Mkt -0.0139 0.0237**
(0.0094) (0.0099)
Contrarian 0.0141 0.0121
(0.0166) (0.0166)
Post Event $\times$ Contrarian -0.0169 -0.0499*
(0.0249) (0.0274)
Momentum -0.0026*** -0.0057***
(0.0007) (0.0018)
Post Event $\times$ Momentum 0.1710** 0.1249
(0.0836) (0.0822)
R-squared 0.0020 0.0111 0.0119 0.0082 0.0021 0.0029 0.0181
R-squared Adj. 0.0020 0.0111 0.0119 0.0082 0.0021 0.0029 0.0180
Observations 119,639 119,639 119,639 119,639 119,639 119,639 119,639
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.2. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.17: OLS Regression Results (5-min interval): Biden Drop Out (Trading Frequency)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Freq Trump Holdings Single Market Trader Types Full
Post Event 0.0061*** 0.0008*** 0.0010*** 0.0014*** 0.0105*** 0.0059*** -0.0053**
(0.0005) (0.0002) (0.0002) (0.0003) (0.0022) (0.0004) (0.0021)
Trad Vol High 0.0012*** 0.0007**
(0.0004) (0.0003)
Post Event $\times$ Trad Vol High 0.0162*** 0.0080***
(0.0013) (0.0011)
Trad Freq High 0.0012** 0.0007
(0.0005) (0.0005)
Post Event $\times$ Trad Freq High 0.0185*** 0.0112***
(0.0015) (0.0014)
Neg Trump Win 0.0007* 0.0005
(0.0004) (0.0004)
Post Event $\times$ Neg Trump Win 0.0154*** 0.0121***
(0.0014) (0.0012)
Single Mkt -0.0004 -0.0001
(0.0004) (0.0005)
Post Event $\times$ Single Mkt -0.0050** 0.0021
(0.0023) (0.0022)
Contrarian 0.0015 0.0011
(0.0020) (0.0020)
Post Event $\times$ Contrarian -0.0026 -0.0083**
(0.0036) (0.0041)
Momentum -0.0005*** -0.0011***
(0.0001) (0.0004)
Post Event $\times$ Momentum 0.0330* 0.0245
(0.0179) (0.0176)
R-squared 0.0021 0.0101 0.0114 0.0087 0.0024 0.0030 0.0175
R-squared Adj. 0.0021 0.0101 0.0114 0.0087 0.0024 0.0030 0.0174
Observations 119,639 119,639 119,639 119,639 119,639 119,639 119,639
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.3. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Table D.18: OLS Regression Results (5-min interval): Biden Drop Out (Trading Intensity)
(1) (2) (3) (4) (5) (6) (7)
Baseline Trading Vol Trading Intensity Trump Holdings Single Market Trader Types Full
Post Event 0.0058*** 0.0008*** 0.0011*** 0.0014*** 0.0095*** 0.0056*** -0.0052***
(0.0004) (0.0002) (0.0002) (0.0002) (0.0017) (0.0004) (0.0017)
Trad Vol High 0.0012*** 0.0006**
(0.0004) (0.0002)
Post Event $\times$ Trad Vol High 0.0154*** 0.0067***
(0.0011) (0.0008)
Trad Int Multi 0.0016*** 0.0011**
(0.0006) (0.0005)
Post Event $\times$ Trad Int Multi 0.0213*** 0.0149***
(0.0015) (0.0013)
Neg Trump Win 0.0008* 0.0004
(0.0004) (0.0004)
Post Event $\times$ Neg Trump Win 0.0146*** 0.0108***
(0.0011) (0.0010)
Single Mkt -0.0006 -0.0002
(0.0005) (0.0006)
Post Event $\times$ Single Mkt -0.0041** 0.0024
(0.0018) (0.0017)
Contrarian 0.0018 0.0012
(0.0023) (0.0023)
Post Event $\times$ Contrarian -0.0040* -0.0117***
(0.0020) (0.0027)
Momentum -0.0005*** -0.0014***
(0.0001) (0.0005)
Post Event $\times$ Momentum 0.0296** 0.0188
(0.0148) (0.0145)
R-squared 0.0023 0.0113 0.0157 0.0097 0.0026 0.0032 0.0215
R-squared Adj. 0.0023 0.0112 0.0157 0.0096 0.0026 0.0032 0.0214
Observations 119,639 119,639 119,639 119,639 119,639 119,639 119,639
  • Notes: This table reports event-specific panel regressions following the same setting in Table C.4. t-statistics use standard errors clustered by trader. * p<<0.1, ** p<<0.05, *** p<<0.01.
Appendix E Additional Event-Time Price Adjustment Plots
This appendix reports supplementary price adjustment plots referenced in Section 5. Unless otherwise noted, all figures use 5-minute event time in a 2-hour pre / 4-hour post window around each event.
(a) Biden YES Token
(b) Biden NO Tokens
Figure E.1: Price and Transaction Volume of Biden Tokens Around Events
Notes: In this plot, each panel plots the 5-minute volume-weighted average transaction price (line) and trading volume (bars) in the event window (-2h, +4h).
(a) Harris YES Token
(b) Harris NO Tokens
Figure E.2: Price and Transaction Volume of Harris Tokens Around Events
Notes: In this plot, each panel plots the 5-minute volume-weighted average transaction price (line) and trading volume (bars) in the event window (-2h, +4h).